# Test time training enhances in-context learning of nonlinear functions

### v1

@@ Abstract
Test-time training (TTT) enhances model performance by explicitly updating designated parameters prior to each prediction to adapt to the test data. While TTT has demonstrated considerable empirical success, its theoretical underpinnings remain limited, particularly for nonlinear models. In this paper, we investigate the combination of TTT with in-context learning (ICL), where the model is given a few examples from the target distribution at inference time. We analyze this framework in the setting of single-index models $y=\sigma_{*}(\left\langle\beta,{\bm{x}}\right\rangle)$, where the feature vector $\beta$ is drawn from a hidden low-dimensional subspace. For single-layer transformers trained with gradient-based algorithms and adopting TTT, we establish an upper bound on the prediction risk. Our theory reveals that TTT enables the single-layer transformers to adapt to both the feature vector $\beta$ and the link function $\sigma_{*}$, which vary across tasks. This creates a sharp contrast with ICL alone, which is theoretically difficult to adapt to shifts in the link function. Moreover, we provide the convergence rate with respect to the data length, showing the predictive error can be driven arbitrarily close to the noise level as the context size and the network width grow.
@@ 1 Introduction
In-context learning (ICL) is a powerful capability of pretrained transformers to solve tasks using a few labeled examples provided as input, without updating their weights. This ability has gained increasing attention with the advent of models with massive context windows, as more examples lead to significantly improved performance (Agarwal et al., 2024). This approach has also proven effective for multimodal tasks (Jiang et al., 2024). From a theoretical perspective, transformers are known to implement algorithms such as linear regression. Recent studies have extended this understanding to nonlinear settings, showing that transformers can learn nonlinear single-index models (Oko et al., 2024) and that softmax attention facilitates data-efficient feature learning (Nishikawa et al., 2025). Nevertheless, ICL faces its inherent limitations: the performance of ICL is fundamentally constrained by factors like the pretraining data (Bigoulaeva et al., 2025) and model architecture (Naim & Asher, 2025).
Test-time training (TTT) has emerged as a promising strategy to overcome these barriers. TTT adapts the model by updating its parameters on the test data before each prediction. This adaptive mechanism has led to strong empirical success across various fields, including large language models (Hu et al., 2025) and video object segmentation (Bertrand et al., 2023). In the context of ICL, TTT can be seamlessly integrated by using the in-context examples as data for task-specific adaptation. For instance, Akyürek et al. (2025) demonstrated that this combination achieves notable improvements on few-shot reasoning benchmarks.
@@ 1.1 Our contribution
${\bm{x}}_{1}^{t},\dots,{\bm{x}}_{N}^{t},{\bm{x}}^{t}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{I}}_{d}),~y_{i}^{t}\approx\sigma_{*}^{t}(\left\langle\beta^{t},{\bm{x}}_{i}^{t}\right\rangle)~(i=1,\cdots,n)$
@@
Theorem 1 (Informal).
Consider learning single-index polynomial $\sigma_{*}^{t}(\left\langle\beta^{t},{\bm{x}}\right\rangle)$ with the transformer trained via Algorithm 1. Then, with probability at least 0.99, we can construct the model $f_{\mathrm{TF}}$ for each prompt that satisfies $\mathbb{E}[|f_{\mathrm{TF}}({\bm{x}})-y|]=\tau+\tilde{O}(m^{-1/2})+\tilde{O}(\sqrt{\frac{r\sqrt{r}}{N_{test}}})$, where $N_{test}$ is the context length in test-time, $m$ is the network width and $\tau=O(1)$ is the noise level, if $N_{pt}=T_{pt}=(d^{\Omega(\mathrm{ie}(\sigma_{*}^{t}))}r^{2})$ and $N_{test}=\tilde{\Omega}(r^{\Omega(\mathrm{ge}(\sigma_{*}^{\mathrm{test}}))})$, where $\mathrm{ie}(\sigma_{*})$ and $\mathrm{ge}(\sigma_{*})$ are the information exponent and the general exponent of the polynomial $\sigma_{*}$, respectively.
@@
  • Efficient sample complexity: Theorem 1 implies that $N_{test}=\tilde{\Omega}(r^{2\vee\Theta(\mathrm{ge}(\sigma_{*}^{\mathrm{test}}))})$ to ensure low predictive loss $o(1)$, which does not depend on the entire dimension $d$. This shows that transformers can adapt to the low-dimensionality of $\beta$. In addition, this statistical complexity does not depend on either the degree of the polynomial $\mathrm{deg}(\sigma_{*}^{\mathrm{test}})$ or $\mathrm{ie}(\sigma_{*}^{\mathrm{test}})$, which outperforms CSQ learners and shows a comparable performance with that of SQ learners.
@@
  • Statistical guarantee for practical settings: We provide an explicit convergence rate with respect to the context length $N_{test}$. This result offers a statistical guarantee in practical scenarios where the dimensions $d,r$ are large but fixed.
@@ 1.2 Related works
In-context learning (Brown et al., 2020) is transformer’s ability to adapt to the specific task using few labeled examples, without updating any parameters. Agarwal et al. (2024) demonstrated that many in-context examples lead to considerably improved performance, and Jiang et al. (2024) confirmed that many-shot ICL is also beneficial for multimodal tasks. The theoretical background of ICL is extensively studied. For example, a wide array of works (Garg et al., 2023; von Oswald et al., 2023; Zhang et al., 2024; Gatmiry et al., 2024) have shown that linear transformers can be trained to perform linear regression in-context. As for nonlinear transformers, Cheng et al. (2024) demonstrated that nonlinear transformers learn to perform gradient descent and thus learn nonlinear functions. Also, Nichani et al. (2024) analyzed learning of causal structure by softmax transformer. Recently, Dherin et al. (2025) showed that ICL in a single-transformer block(a self-attention layer and subsequent MLP layer) corresponds to low-rank update in MLP layer. Regarding limitations of ICL, Bigoulaeva et al. (2025) argued that pretraining datasets impose a fundamental limit on the model’s capability with ICL. Furthermore, (Naim & Asher, 2025) demonstrated that the transformer’s ICL ability to generalize functions is limited to certain input values, and found that this limitation comes from layer normalization and softmax attention.
@@
Test-time training (Sun et al. (2020), Liu et al. (2021)) updates the model using test data before making predictions, thereby addressing distribution shifts. TTT achieved success in many fields. For example, Bertrand et al. (2023) applied TTT for the video object segmentation task and achieved a significant improvement in the performance. Moreover, Hu et al. (2025) analyzed test-time learning of large language models and achieved at least 20% higher performance on domain knowledge adaptation. Furthermore, Zhang et al. (2025) proposed adopting a large chunk update, and validated the effectiveness of their approach to long-context data through tasks like image sets and language model. As for TTT combined with few-shot prediction, Akyürek et al. (2025) reported that introducing TTT with in-context examples resulted in 6 times higher accuracy in the Abstraction and Reasoning Corpus and a 7.3 percent higher score on BIG-Bench Hard. Finally, regarding the theory behind TTT for in-context learning, Gozeten et al. (2025) analyzed the linear transformer with a single gradient step and characterized the prediction risk of the model with TTT, showing that TTT can mitigate distribution shift.
@@ 2 Preliminaries and problem settings
Let $\mathrm{He}_{i}(z)=(-1)^{i}\mathrm{e}^{\frac{z^{2}}{2}}\frac{\mathrm{d}^{i}}{\mathrm{d}z^{i}}\mathrm{e}^{\frac{-z^{2}}{2}}$ be the degree-$i$ (probabilist’s) Hermite polynomial. $\mathbb{S}^{d-1}$ denotes the unit sphere in $\mathbb{R}^{d}$. For matrix $\bm{A}$}, we denote its $\ell_{2}$ operator norm and Frobenius norm as $\|{\bm{A}}\|_{2}$ and $\|{\bm{A}}\|_{F}$, respectively. For a set $\mathrm{S}$, $\mathrm{Unif(S)}$ denotes the uniform distribution over $\mathrm{S}$. $\tilde{O},\tilde{\Omega},\tilde{\Theta}$ means $O,\Omega,\Theta$ where polylogamistic terms of $d$ and $1/\varepsilon$ are hidden. $O_{d},\Omega_{d},\Theta_{d}$ means the order with respect to the dimension $d,r$, while $\Theta_{\varepsilon}$ denotes the order in terms of $\varepsilon$.
@@ 2.1 In-context learning and test-time training
We consider the basic setting in ICL, which is introduced by Garg et al. (2023) (see Oko et al. (2024) and Lee et al. (2024)). In this setting, the model is given a sequence $({\bm{x}}_{1},y_{1},\cdots,{\bm{x}}_{N},y_{N},{\bm{x}})$ called prompt. The labeled pairs $({\bm{x}}_{i},y_{i})$ are called contexts, and then the model is asked to predict the output that corresponds to $\bm{x}$}. The context is sometimes abbreviated as $({\bm{X}}_{n},{\bm{y}}_{n})$ where ${\bm{X}}_{n}=({\bm{x}}_{1},\cdots,{\bm{x}}_{n}),{\bm{y}}_{n}=(y_{1},\cdots,y_{n})$. In this work, we assume that the $\bm{x}$} and $y$ are generated as follows:
${\bm{x}}_{1}^{t},\dots,{\bm{x}}_{N}^{t},{\bm{x}}^{t}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{I}}_{d}),~y_{i}^{t}=f_{*}^{t}({\bm{x}}_{i}^{t})+\zeta_{i},~\zeta_{i}\sim\mathrm{Unif}({-\tau,\tau}).$
ICL aims to predict $y=f_{*}^{t}({\bm{x}})+\zeta$ by mere observation of the context, without updating parameters for each prompt. However, we introduce test-time training to further enhance the model’s accuracy. In pretraining, we train the model with $T_{pt}$ distinct datasets $({\bm{x}}_{1}^{t},y_{1}^{t},\cdots,{\bm{x}}_{N_{pt}}^{t},y_{N_{pt}}^{t},{\bm{x}}^{t},y^{t})_{t=1}^{T_{pt}}$. In test-time, we divide the context into four groups with $i$-th group’s length $N_{i}$. This means the test-time prompt is $({\bm{X}}_{N_{1}},{\bm{y}}_{N_{1}},{\bm{X}}_{N_{2}},{\bm{y}}_{N_{2}},{\bm{X}}_{N_{3}},{\bm{y}}_{N_{3}},{\bm{X}}_{N_{4}},{\bm{y}}_{N_{4}},{\bm{x}})$. Each group of data plays a different role in test-time training: See Subsection 2.4 for the detail.
@@ 2.2 Single index model
We consider single-index models for the input-output relationship, where the output depends solely on the direction of the feature vector $\beta$. To predict this relationship accurately, the models are expected to learn the target direction $\beta$ from the high-dimensional data in $\mathbb{R}^{d}$. Consequently, single-index models have been extensively studied in machine learning theory (Bai & Lee, 2020; Ba et al., 2022; Bietti et al., 2022; Mousavi-Hosseini et al., 2023; Berthier et al., 2024), particularly to examine adaptability to the low-dimensional subspace. The target function $f_{*}^{t}({\bm{x}})$ is determined as:
$f_{*}^{t}({\bm{x}})=\sigma_{*}^{t}(\left\langle\beta^{t},{\bm{x}}\right\rangle).$
@@
    $\mathbb{E}[c_{Q}^{t}]=\Theta(1)\neq 0,~\sum_{i=Q}^{P}{(c_{i}^{t})}^{2}\leq R_{c}=\Theta(1)~(\text{a.s.})~\text{and}~(c_{Q}^{t},\cdots,c_{P}^{t})\neq(0,\cdots,0)~(\text{a.s.}).$
Remark 2.
@@
  • The information exponent (Dudeja & Hsu, 2018; Arous et al., 2021) $\mathrm{ie}(\sigma_{*})$ is the smallest non-zero degree of the hermite expansion of $\sigma_{*}$.
@@ 2.3 Student model
We employ a single-layer transformer model for the student model. In pretraining, we use the same model as Nishikawa et al. (2025). We first construct the embedding $\bm{E}$} as
${\bm{E}}=\begin{bmatrix}{\bm{x}}_{1}&\cdots&{\bm{x}}_{N}&{\bm{x}}\\ y_{1}&\cdots&y_{N}&1\end{bmatrix}\in\mathbb{R}^{(d+1)\times(N+1)}$. Then, we apply the softmax attention layer as
$\mathrm{Attn}({\bm{E}})={\bm{W}}^{V}{\bm{E}}\cdot\mathrm{softmax}(\mathrm{Mask}(\rho^{-1}\cdot({\bm{W}}^{K}{\bm{E}})^{\top}{\bm{W}}^{Q}{\bm{E}})),$
@@
$f_{\mathrm{IC}}({\bm{x}})$ $=\mathrm{MLP}\circ\mathrm{Attn}({\bm{E}})_{:,N+1}={\bm{a}}^{\top}\sigma({\bm{W}}^{F}\mathrm{Attn}({\bm{E}})_{:,N+1}+{\bm{b}}),$
where $\sigma$ is applied entry-wise. Finally, we adopt some simplification. Let ${\bm{W}}^{KQ}=({\bm{W}}^{K})^{\top}{\bm{W}}^{Q}\in\mathbb{R}^{(d+1)\times(d+1)}$ and ${\bm{W}}^{FV}={\bm{W}}^{F}{\bm{W}}^{V}\mathbb{R}^{(m+1)\times(d+1)}$. We use the following parametrization:
@@
for $\Gamma\in\mathbb{R}^{d\times d}$ and ${\bm{v}}\in\mathbb{R}^{(d+1)\times 1}$. This kind of simplification, specifying some of the parameters as zero, is often adopted in many theoretical works on transformers (Zhang et al., 2023; Huang et al., 2023; Kim & Suzuki, 2025). Overall, the model’s output is written as
$f_{\mathrm{IC}}({\bm{x}})=\sum_{j=1}^{m}a_{j}\sigma\left(v_{j}\frac{\sum_{i=1}^{N}y_{i}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{\top}\Gamma{\bm{x}}/\rho}}{\sum_{i=1}^{N}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{\top}\Gamma{\bm{x}}/\rho}}+b_{j}\right).$
@@
Remark 4.
The use of in-context data in attention output leads to data-efficient prediction, but it prevents the model from achieving near-zero error. This is because of the high order term of $\left\langle\beta,{\bm{x}}\right\rangle$. When $N$ is sufficiently large, the denominator of the attention output $\sum_{i=1}^{N}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{\top}\Gamma{\bm{x}}/\rho}$ approximates $\mathbb{E}[\exp\sigma_{*}({\left\langle\beta,{\bm{x}}_{1}\right\rangle/\rho})\exp(\left\langle\Gamma_{*}{\bm{x}},{\bm{x}}_{1}\right\rangle/\rho)]$, as explained in Nishikawa et al. (2025). The key of their work is, since $\exp(\left\langle\Gamma_{*}{\bm{x}},{\bm{x}}_{1}\right\rangle/\rho)$ contains all the Hermite coefficients and $\exp\sigma_{*}({\left\langle\beta,{\bm{x}}_{1}\right\rangle/\rho})$ has nonzero coefficient for $\mathrm{He}_{\mathrm{ge}(\sigma_{*})}$, this attention layer can compute $\left\langle\beta,{\bm{x}}\right\rangle^{\mathrm{ge}(\sigma_{*})}$. However, this means that the output of the attention module contains $\left\langle\beta,{\bm{x}}\right\rangle^{i}$ for $i\geq\mathrm{ge}(\sigma_{*})+1$ as well. This causes inevitable prediction error, even when $N\rightarrow\infty$. Meanwhile, when we find $\hat{{\bm{u}}}$ that is sufficiently near the feature vector $\beta$, we can achieve near-zero error. Therefore we do not use in-context data in the final prediction $f_{\mathrm{TF}}$.

### v2

@@ Abstract
Test-time training (TTT) enhances model performance by explicitly updating designated parameters prior to each prediction to adapt to the test data. While TTT has demonstrated considerable empirical success, its theoretical underpinnings remain limited, particularly for nonlinear models. In this paper, we investigate the combination of TTT with in-context learning (ICL), where the model is given a few examples from the target distribution at inference time. We analyze this framework in the setting of single-index models $y=\sigma_{*}(\left\langle\beta,\bf{x}\right\rangle)$, where the feature vector $\beta$ is drawn from a hidden low-dimensional subspace. For single-layer transformers trained with gradient-based algorithms and adopting TTT, we establish an upper bound on the prediction risk. Our theory reveals that TTT enables the single-layer transformers to adapt to both the feature vector $\beta$ and the link function $\sigma_{*}$, which vary across tasks. This creates a sharp contrast with ICL alone, which is theoretically difficult to adapt to shifts in the link function. Moreover, we provide the convergence rate with respect to the data length, showing the predictive error can be driven arbitrarily close to the noise level as the context size and the network width grow.
Machine Learning, ICML
@@ 1 Introduction
In-context learning (ICL) is a powerful capability of pretrained transformers to solve tasks using a few labeled examples provided as input, without updating their weights. This ability has gained increasing attention with the advent of models with massive context windows, as more examples lead to significantly improved performance (Agarwal et al., 2024). This approach has also proven effective for multimodal tasks (Jiang et al., 2024). From a theoretical perspective, transformers are known to implement algorithms such as linear regression. Recent studies have extended this understanding to nonlinear settings, showing that transformers can learn nonlinear single-index models (Oko et al., 2024) and that softmax attention facilitates data-efficient feature learning (Nishikawa et al., 2025). Nevertheless, ICL faces its inherent limitations: the performance of ICL is fundamentally constrained by factors like the pretraining data (Bigoulaeva et al., 2025) and model architecture (Naim and Asher, 2025).
Test-time training (TTT) has emerged as a promising strategy to overcome these barriers. TTT adapts the model by updating its parameters on the test data before each prediction. This adaptive mechanism has led to strong empirical success across various fields, including large language models (Hu et al., 2025) and video object segmentation (Bertrand et al., 2023). In the context of ICL, TTT can be seamlessly integrated by using the in-context examples as data for task-specific adaptation. For instance, akyürek2025surprisingeffectivenesstesttimetraining demonstrated that this combination achieves notable improvements on few-shot reasoning benchmarks.
@@ 1.1 Our contribution
${\bm{x}}_{1}^{t},\dots,{\bm{x}}_{N}^{t},{\bm{x}}^{t}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{I}}_{d}),~y_{i}^{t}\approx\sigma_{*}^{t}(\left\langle\beta^{t},{\bm{x}}_{i}^{t}\right\rangle)$~
$(i=1,\cdots,n)$
@@
Theorem 1.1 (Informal).
Consider learning single-index polynomial $\sigma_{*}^{t}(\left\langle\beta^{t},{\bm{x}}\right\rangle)$ with the transformer trained via Algorithm 1. Then, with probability at least 0.99, we can construct the model $f_{\mathrm{TF}}$ for each prompt that satisfies $\mathbb{E}[|f_{\mathrm{TF}}({\bm{x}}^{t})-\sigma_{*}^{t}(\left\langle\beta^{t},{\bm{x}}^{t}\right\rangle)|]=\tilde{O}(m^{-1/2})+\tilde{O}(\sqrt{\frac{r\sqrt{r}}{N_{test}}})$, where $N_{pt}$ and $T_{pt}$ are the context length and the number of tasks in pretraining, respectively, $N_{test}$ is the context length in test-time, $m$ is the network width, if $N_{pt}=T_{pt}=(d^{\Omega(\mathrm{ie}(\sigma_{*}^{t}))}r^{2})$ and $N_{test}=\tilde{\Omega}(r^{\Omega(\mathrm{ge}(\sigma_{*}^{\mathrm{test}}))})$, where $\mathrm{ie}(\sigma_{*})$ and $\mathrm{ge}(\sigma_{*})$ are the information exponent and the general exponent of the polynomial $\sigma_{*}$, respectively.
@@
  • Efficient sample complexity: Theorem 1.1 implies that $N_{test}=\tilde{\Omega}(r^{2\vee\Theta(\mathrm{ge}(\sigma_{*}^{\mathrm{test}}))})$ to ensure low predictive loss $\tilde{O}(\sqrt{\frac{r\sqrt{r}}{N_{test}}})$, which does not depend on the entire dimension $d$. This shows that transformers can adapt to the low-dimensionality of $\beta$. In addition, this statistical complexity does not depend on either the degree of the polynomial $\mathrm{deg}(\sigma_{*}^{\mathrm{test}})$ or $\mathrm{ie}(\sigma_{*}^{\mathrm{test}})$, which outperforms CSQ learners and shows a comparable performance with that of SQ learners.
@@
  • Statistical guarantee for practical settings: We provide an explicit convergence rate with respect to the context length $N_{test}$. This result offers a statistical guarantee in practical scenarios where the dimensions $d,r$ are large but fixed. This establishes a fundamental advantage over standard ICL: while existing frameworks (Nishikawa et al., 2025) fail to prove that the predictive risk vanishes as $N_{test}\to\infty$, our TTT-based approach ensures that the error can be driven to an arbitrarily small value.
To empirically validate our theoretical predictions, we perform numerical experiments with a 2-layer GPT-2 model in a controlled setting. The results, presented in Figure 1, reveal that while standard ICL struggles to adapt to shifts in the link function $\sigma_{*}$, TTT achieves significantly lower predictive error, which continues to decrease with increasing sequence length. This empirical success validates our theoretical claim: TTT provides the necessary flexibility to achieve exact recovery of task-specific nonlinearities, beyond the architectural limitations of static in-context learning. See section 4 for further details.
@@ 1.2 Related works
In-context learning (Brown et al., 2020) is transformer’s ability to adapt to the specific task using few labeled examples, without updating any parameters. Agarwal et al. (2024) demonstrated that many in-context examples lead to considerably improved performance, and Jiang et al. (2024) confirmed that many-shot ICL is also beneficial for multimodal tasks. The theoretical background of ICL is extensively studied. For example, a wide array of works (Garg et al., 2023; von Oswald et al., 2023; Zhang et al., 2024b; Gatmiry et al., 2024) have shown that linear transformers can be trained to perform linear regression in-context. As for nonlinear transformers, Cheng et al. (2024) demonstrated that nonlinear transformers learn to perform gradient descent and thus learn nonlinear functions. Also, Nichani et al. (2024) analyzed learning of causal structure by softmax transformer. Recently, Dherin et al. (2025) showed that ICL in a single-transformer block (a self-attention layer and subsequent MLP layer) corresponds to low-rank update in MLP layer. Regarding limitations of ICL, Bigoulaeva et al. (2025) argued that pretraining datasets impose a fundamental limit on the model’s capability with ICL. Furthermore, (Naim and Asher, 2025) demonstrated that the transformer’s ICL ability to generalize functions is limited to certain input values, and found that this limitation comes from layer normalization and softmax attention.
@@
Test-time training (Sun et al., 2020; Liu et al., 2021) updates the model using test data before making predictions, thereby addressing distribution shifts. TTT achieved success in many fields. For example, Bertrand et al. (2023) applied TTT for the video object segmentation task and achieved a significant improvement in the performance. Moreover, Hu et al. (2025) analyzed test-time learning of large language models and achieved at least 20% higher performance on domain knowledge adaptation. Furthermore, Zhang et al. (2025) proposed adopting a large chunk update, and validated the effectiveness of their approach to long-context data through tasks like image sets and language model. As for TTT combined with few-shot prediction, akyürek2025surprisingeffectivenesstesttimetraining reported that introducing TTT with in-context examples resulted in 6 times higher accuracy in the Abstraction and Reasoning Corpus and a 7.3 percent higher score on BIG-Bench Hard. Finally, regarding the theory behind TTT for in-context learning, Gozeten et al. (2025) analyzed the linear transformer with a single gradient step and characterized the prediction risk of the model with TTT, showing that TTT can mitigate distribution shift.
@@ 2 Preliminaries and problem settings
Let $\mathrm{He}_{i}(z)=(-1)^{i}\mathrm{e}^{\frac{z^{2}}{2}}\frac{\mathrm{d}^{i}}{\mathrm{d}z^{i}}\mathrm{e}^{\frac{-z^{2}}{2}}$ be the degree-$i$ (probabilist’s) Hermite polynomial. $\mathbb{S}^{d-1}$ denotes the unit sphere in $\mathbb{R}^{d}$. For matrix $\bm{A}$}, we denote its $\ell_{2}$ operator norm and Frobenius norm as $\|{\bm{A}}\|_{2}$ and $\|{\bm{A}}\|_{F}$, respectively. For a set $\mathrm{S}$, $\mathrm{Unif(S)}$ denotes the uniform distribution over $\mathrm{S}$. $\tilde{O},\tilde{\Omega},\tilde{\Theta}$ means $O,\Omega,\Theta$ where polylogarithmic terms of $d$ and $1/\varepsilon$ are hidden. $O_{d},\Omega_{d},\Theta_{d}$ means the order with respect to the dimension $d,r$, while $\Theta_{\varepsilon}$ denotes the order in terms of $\varepsilon$.
@@ 2.1 In-context learning and test-time training
We consider the basic setting in ICL, which is introduced by Garg et al. (2023) (see Oko et al. (2024) and Lee et al. (2024)). In this setting, the model is given a sequence $({\bm{x}}_{1},y_{1},\cdots,{\bm{x}}_{N},y_{N},{\bm{x}})$ called prompt. The labeled pairs $({\bm{x}}_{i},y_{i})\in\mathbb{R}^{d}\times\mathbb{R}$ are called contexts, and ${\bm{x}}\in\mathbb{R}^{d}$ is referred to as query. The model is asked to predict the output that corresponds to $\bm{x}$} based on the context. The context is sometimes abbreviated as $({\bm{X}}_{n},{\bm{y}}_{n})$ where ${\bm{X}}_{n}=({\bm{x}}_{1},\cdots,{\bm{x}}_{n}),{\bm{y}}_{n}=(y_{1},\cdots,y_{n})$. In this work, we assume that the $\bm{x}$} and $y$ are generated as follows:
${\bm{x}}_{1}^{t},\dots,{\bm{x}}_{N}^{t},{\bm{x}}^{t}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{I}}_{d}),~y_{i}^{t}=f_{*}^{t}({\bm{x}}_{i}^{t})+\zeta_{i},$
$~\zeta_{i}\sim\mathrm{Unif}({-\tau,\tau}).$
ICL aims to predict $y=f_{*}^{t}({\bm{x}})+\zeta$ by mere observation of the context, without updating parameters for each prompt. However, we introduce test-time training to further enhance the model’s accuracy. In pretraining, we train the model with $T_{pt}$ distinct datasets $({\bm{x}}_{1}^{t},y_{1}^{t},\cdots,{\bm{x}}_{N_{pt}}^{t},y_{N_{pt}}^{t},{\bm{x}}^{t},y^{t})_{t=1}^{T_{pt}}$, with each prompts consisting of $N_{pt}$ queries. In test-time, we divide the context into four groups with $i$-th group’s length $N_{i}$. This means the test-time prompt is $({\bm{X}}_{N_{1}},{\bm{y}}_{N_{1}},{\bm{X}}_{N_{2}},{\bm{y}}_{N_{2}},{\bm{X}}_{N_{3}},{\bm{y}}_{N_{3}},{\bm{X}}_{N_{4}},{\bm{y}}_{N_{4}},{\bm{x}})$. Each group of data plays a different role in test-time training: See subsection 2.4 for the detail. Let $N_{test}=\sum_{i=1}^{4}N_{i}$ be the total number of contexts test-time.
@@ 2.2 Single index model
We consider single-index models for the input-output relationship, where the output depends solely on the direction of the feature vector $\beta$. To predict this relationship accurately, the models are expected to learn the target direction $\beta$ from the high-dimensional data in $\mathbb{R}^{d}$. Consequently, single-index models have been extensively studied in machine learning theory (Bai and Lee, 2020; Ba et al., 2022; Bietti et al., 2022; Mousavi-Hosseini et al., 2023; Berthier et al., 2024), particularly to examine adaptability to the low-dimensional subspace. The target function $f_{*}^{t}({\bm{x}})$ is determined as $f_{*}^{t}({\bm{x}})=\sigma_{*}^{t}(\left\langle\beta^{t},{\bm{x}}\right\rangle).$
@@
    $\mathbb{E}[c_{Q}^{t}]=\Theta(1)\neq 0,~\sum_{i=Q}^{P}{(c_{i}^{t})}^{2}\leq R_{c}=\Theta(1)~(\text{a.s.})$~
    $\text{and}~(c_{Q}^{t},\cdots,c_{P}^{t})\neq(0,\cdots,0)~(\text{a.s.}).$
Remark 2.1.
@@
  • The information exponent (Dudeja and Hsu, 2018; Arous et al., 2021) $\mathrm{ie}(\sigma_{*})$ is the smallest non-zero degree of the hermite expansion of $\sigma_{*}$.
@@ 2.3 Student model
To facilitate our theoretical analysis, we need to establish a concrete architecture for the student model. We employ a single-layer transformer model, formulated as follows. In pretraining, we use the same model as Nishikawa et al. (2025). We first construct the embedding $\bm{E}$} as
${\bm{E}}=\begin{bmatrix}{\bm{x}}_{1}&\cdots&{\bm{x}}_{N}&{\bm{x}}\\ y_{1}&\cdots&y_{N}&1\end{bmatrix}\in\mathbb{R}^{(d+1)\times(N+1)}.$
Then, we apply the softmax attention layer as
$\mathrm{Attn}({\bm{E}})$
$=$ ${\bm{W}}^{V}{\bm{E}}\cdot\mathrm{softmax}(\mathrm{Mask}(\rho^{-1}\cdot({\bm{W}}^{K}{\bm{E}})^{\top}{\bm{W}}^{Q}{\bm{E}})),$
@@
$f_{\mathrm{IC}}(\Gamma,{\bm{X}}_{N},{\bm{y}}_{N},{\bm{x}})$ $=\mathrm{MLP}\circ\mathrm{Attn}({\bm{E}})_{:,N+1}$
$={\bm{a}}^{\top}\sigma({\bm{W}}^{F}\mathrm{Attn}({\bm{E}})_{:,N+1}+{\bm{b}}),$
where $\sigma$ is applied entry-wise. Finally, we adopt some simplification. Let ${\bm{W}}^{KQ}=({\bm{W}}^{K})^{\top}{\bm{W}}^{Q}\in\mathbb{R}^{(d+1)\times(d+1)}$ and ${\bm{W}}^{FV}={\bm{W}}^{F}{\bm{W}}^{V}\in\mathbb{R}^{(m+1)\times(d+1)}$. We use the following parametrization:
@@
for $\Gamma\in\mathbb{R}^{d\times d}$ and ${\bm{v}}\in\mathbb{R}^{(d+1)\times 1}$. This kind of simplification, specifying some of the parameters as zero, is often adopted in many theoretical works on transformers (Zhang et al., 2024a; Huang et al., 2023; Kim and Suzuki, 2025). Overall, the model’s output is written as
$f_{\mathrm{IC}}(\Gamma,{\bm{X}}_{N},{\bm{y}}_{N},{\bm{x}})$
$=$ $\sum_{j=1}^{m}a_{j}\sigma\left(v_{j}\frac{\sum_{i=1}^{N}y_{i}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{\top}\Gamma{\bm{x}}/\rho}}{\sum_{i=1}^{N}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{\top}\Gamma{\bm{x}}/\rho}}+b_{j}\right).$
@@
Remark 2.3.
Incorporating in-context data into the final prediction—a standard practice in existing TTT frameworks (akyürek2025surprisingeffectivenesstesttimetraining)—presents a fundamental trade-off between finite-sample efficiency and asymptotic consistency. In the small-data regime (small $N$), retaining in-context examples is beneficial, as it compensates for the estimation error of the updated parameters. However, this benefit comes at the cost of an architectural bottleneck in the large-data regime ($N\to\infty$): the use of the attention mechanism adds unavoidable noise. As Nishikawa et al. (2025) explains, when the context length $N$ is sufficiently large, the denominator of the attention output $N^{-1}\sum_{i=1}^{N}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{\top}\Gamma{\bm{x}}/\rho}$ approximates $\mathbb{E}[\exp\sigma_{*}({\left\langle\beta,{\bm{x}}_{1}\right\rangle/\rho})\exp(\left\langle\Gamma_{*}{\bm{x}},{\bm{x}}_{1}\right\rangle/\rho)]$. The key of their work is, since $\exp(\left\langle\Gamma_{*}{\bm{x}},{\bm{x}}_{1}\right\rangle/\rho)$ contains all the Hermite coefficients and $\exp\sigma_{*}({\left\langle\beta,{\bm{x}}_{1}\right\rangle/\rho})$ has nonzero coefficient for $\mathrm{He}_{\mathrm{ge}(\sigma_{*})}$, this attention layer can compute $\left\langle\beta,{\bm{x}}\right\rangle^{\mathrm{ge}(\sigma_{*})}$. However, this mechanism inevitably retains $\left\langle\beta,{\bm{x}}\right\rangle^{k}$ for degrees $k>\mathrm{ge}(\sigma_{*})$, causing an unavoidable prediction error (bias) that persists even as $N\rightarrow\infty$. By contrast, our final predictor omits in-context data and relies solely on the learned parameter $\hat{{\bm{u}}}$, thereby avoiding this asymptotic bias. We adopt this design to demonstrate that TTT can surpass the standard ICL limit and attain exact recovery.